OpenStudy (anonymous):

evaulate the difference quotient f(x)=(x+3)/(x+1), f(x)-f(1)/x-1

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

How much is f(1)? If f(x) = (x+3)/(x+1) Then f(1) = (1+3)/(1+1) = 4/2 = 2 So f(x) - f(1)/(x-1) = \(\frac{x+3}{x+1}\) - \(\frac{2}{x-1}\) Understood? :) Now calculate that! :D

OpenStudy (anonymous):

isnt the x-1 should be under the whole thing

OpenStudy (akashdeepdeb):

The question is framed in such a way that it seems like it is under f(1) only. :D

OpenStudy (akashdeepdeb):

Then calculate \[\frac { \frac{ x+3 }{ x+1 } - 2 }{x-1}\]

OpenStudy (anonymous):

that is the part i am having problem with

OpenStudy (anonymous):

forgot how to rationalize it

OpenStudy (akashdeepdeb):

It is already rational no need to rationalize.. \[\frac{ (x+3) - 2(x+1) }{ (x+1)(x-1) }\] Getting this?

OpenStudy (anonymous):

yes thank you! :D

OpenStudy (akashdeepdeb):

:)